### v1

# LLM Security Guard for Code

Abstract.
Many developers rely on Large Language Models (LLMs) to facilitate software development. Nevertheless, these models have exhibited limited capabilities in the security domain. We introduce LLMSecGuard, an open-source framework that offers enhanced code security through the synergy between static code analyzers and LLMs. LLMSecGuard aims to equip practitioners with code solutions that are more secure than the code initially generated by LLMs. It also benchmarks LLMs, providing valuable insights into the evolving security properties of these models.
Security analysis, secure code generation, code models
journalyear: 2024conference: 28th International Conference on Evaluation and Assessment in Software Engineering; June 18–21, 2024; Salerno, Italybooktitle: 28th International Conference on Evaluation and Assessment in Software Engineering (EASE 2024), June 18–21, 2024, Salerno, Italycopyright: nonedoi: 10.1145/xxxccs: Security and privacy Software security engineering
1. Introduction
The use of Large Language Models (LLMs) such as ChatGPT and Copilot has become popular for software development activities such as coding, design, comprehension, etc. (Hou et al., 2023; Fan et al., 2023). Nevertheless, hallucination, i.e., “presenting incorrect information as if it is correct”, poses serious challenges for LLM users (Huang et al., 2023). This issue is more prevalent in domains where reliable training content is scarce, such as in the realm of code security. Indeed, recent studies have shown that code models are widely adopted for code generation (GitHub, 2023), but they have limited capabilities in the software security domain (Wang et al., 2023; Asare et al., 2023). Therefore, a vulnerable code piece that an LLM mistakenly recommends as a secure solution could compromise the entire system’s security if it is used without enough scrutiny.
We introduce LLMSecGuard, a framework designed to extensively examine the security properties of LLMs and leverage their analytical capabilities for secure code development. It applies static security analysis on LLM-generated code to uncover potential security issues and guides LLMs in resolving such issues in the code. Moreover, LLMSecGuard assesses the security properties of LLMs and benchmark them across different CWEs. In summary, LLMSecGuard contributes to more secure software development. Unlike existing code assistants, developers can integrate unlimited LLMs and code analysis engines into this framework through REST APIs. LLMSecGuard is open-source and publicly available on GitHub. [^1] https://github.com/aryakvnust/LLMSecGuard
The remainder of this paper is structured as follows. In Section 2, we motivate this work. In Section 3, we introduce LLMSecGuard, and in Section 4, we explain its two use cases. In Section 5, we present related work. In Section 6, we outline our plans, and in Section 7, we conclude this paper.
2. Motivation
Security issues are pervasive in multiple domains (Bühlmann and Ghafari, 2022), from mobile applications (Ghafari et al., 2017; Gadient et al., 2019) and web servers (Gadient et al., 2021), to critical software systems (Wetzels et al., 2023). There are program analysis tools designed to uncover security issues, but studies indicate their limitations (Esposito et al., 2024; Corrodi et al., 2018), as well as their lack of popularity among developers (Hazhirpasand and Ghafari, 2021). Unfortunately, the security landscape may not improve as we witness the popularity of language models for code generation (GitHub, 2023). It is concerning that developers place undue trust in these models, which are known to generate insecure code examples (Wang et al., 2023; Asare et al., 2023).
In the following, we explain how we could iteratively ask ChatGPT 3.5 to generate secure code examples for symmetric encryption. Encryption plays a crucial role in our digital world, but its correct adoption has proven to be challenging for developers (Hazhirpasand et al., 2019). For instance, the analysis of 489 open-source Java projects showed that 85% of cryptography APIs are misused (Hazhirpasand et al., 2020). Importantly, existing tools that identify and fix such issues do not fulfill developers’ expectations (Afrose et al., 2023; Zhang et al., 2023).
Consider the StackOverflow post ID 3451670, which suffers from security issues such as “ECB encryption mode”, “static or constant key”, “static salt for key derivation”, and “hardcoded password”. When we presented this inquiry to ChatGPT, requesting a “secure Java code example”, the resulting response still contained identical security issues.
We then explicitly instructed ChatGPT that “I expect a secure encryption algorithm, a secure encryption mode, and a secure Initialization Vector (IV). Do not use hard-coded secrets, suggest best practices, and make sure that your example provides confidentiality and integrity”. Listing 1 presents the response. [^2] https://chat.openai.com/share/8ce9eb6e-88b7-44dc-807a-51d04e7d446c
1public class SecureEncryptionExample {
2 public static void main(String[] args) throws Exception {
3 String username = "bob@google.org";
4 String password = "Password1";
5 String secretID = "BlahBlahBlah";
6 String SALT2 = "deliciously salty";
7 // Generate a secure random IV (Initialization Vector)
8 SecureRandom secureRandom = new SecureRandom();
9 byte[] iv = new byte[12];
10 secureRandom.nextBytes(iv);
11 // Derive a key from the password using a secure method
12 byte[] key = deriveKey(SALT2, username, password);
13 // Create a SecretKey object using the derived key
14 SecretKey secretKey = new SecretKeySpec(key, "AES");
15 // Instantiate the cipher with AES in GCM mode
16 Cipher cipher = Cipher.getInstance("AES/GCM/NoPadding");
17 ...
18 }
19 private static byte[] deriveKey(String salt, String username, String password) throws Exception {
20 // Concatenate salt, username, and password
21 String combined = salt + username + password;
22 // Use a secure hash function (SHA-256) to derive a key
23 MessageDigest md = MessageDigest.getInstance("SHA-256");
24 byte[] hash = md.digest(combined.getBytes());
25 return Arrays.copyOf(hash, 16);
26 }
27 }
Listing 1: ChatGPT’s improved response to Post ID 3451670
ChatGPT identified the constant key, and it correctly suggested the secure “GCM” encryption mode as a best practice. However, violations such as “static salt for key generation” and “hardcoded password” were in the response without any proper warning. The suggested code, at line 23, included the “SHA-256” hash function for key generation, which is not a best practice. Anyway, when we asked about the security risk at line 6 (i.e., static salt), ChatGPT responded: “Using a hard-coded string as a salt is generally not considered a good practice for security. Ideally, salt should be unique for each user and should be randomly generated for each password. Using a constant salt for all users could make the system more vulnerable to certain types of attacks, such as rainbow table attacks”. It then provided the code example in Listing 2 as a replacement.
1// Generate a secure random salt
2SecureRandom secureRandom = new SecureRandom();
3byte[] salt = new byte[16];
4secureRandom.nextBytes(salt);
5// Convert the salt to a Base64-encoded string for storage
6String SALT2 = Base64.getEncoder().encodeToString(salt);
Listing 2: ChatGPT’s suggestion to replace a static salt
Subsequently, we asked, “Are there any other security risks in your provided code?”. It failed to detect the hardcoded password, but at this stage, it correctly suggested a key derivation function (KDF) instead of the hash function for generating a password-based key. Upon a direct inquiry about the line where a hardcoded password existed, it offered a secure suggestion.
Through proper interactions with ChatGPT, specifically by listing potential security issues and their locations within the code, it is feasible to harness ChatGPT’s power to generate secure code examples. However, novice developers may not possess the necessary expertise to provide such inputs. Hence, we introduce LLMSecGuard, a framework that links LLMs and static analysis tools to overcome this limitation.
3. LLMSecGuard
We introduce LLMSecGuard, an open-source framework that offers enhanced code security through the synergy between code security analyzers and LLMs. The primary objective of LLMSecGuard is (i) to enhance the security of LLM-generated code, and (ii) to benchmark the security properties of LLMs. It adopts a RESTful architecture, implemented in Python using Django and Flask, and uses SQLite for data persistence. LLMSecGuard is publicly available on GitHub. [^3] https://github.com/aryakvnust/LLMSecGuard
Figure 1. LLMSecGuard’s components
Figure 1 illustrates the schematic architecture of LLMSecGuard, comprising three main components: “Prompt Agent”, “Security Agent”, and “Benchmark Agent”.
Prompt Agent. This component is responsible for receiving a prompt and providing other components with the code model’s response. Upon receipt of a prompt, “Prompt Agent” undertakes the task of engineering a response. Particularly, it can reformulate a prompt, pass the prompt to code models, collect the response, and forward the result to other components.
Security Agent. This component has a pivotal role in uncovering security issues in LLM-generated code. Precisely, it is entrusted to pass the code to static code analysis engines (such as Semgrep and Weggli), and to collect potential security vulnerabilities.
Benchmark Agent. This component puts different LLMs to security test. Particularly, it evaluates the security properties of LLMs based on a set of challenges, structured in JSON format. Each challenge includes a prompt and the expected outcome, and “Benchmark Agent” is responsible for flagging LLMs that pass the challenge.
4. Usage Scenarios
We describe the LLMSecGuard’s main configurations including its connection with external entities such as LLMs and code analyzers. We then explain the interplay between different components through two main usage scenarios, namely “benchmarking” and “code generation”.
4.1. Setup
There are several key configurations for running LLMSecGuard that users can adjust according to their preferences.
LLMSecGuard requires a minimum of one LLM and one code analysis engine to operate effectively. This can be easily set up by providing the API information (e.g., API endpoint and API key) associated with each entity. There is no limitation, and users can add as many entities as they wish. Currently, we have instantiated LLMSecGuard with ANYSCALE Llama2 (Meta AI, date), and Weggli and Semgrep static code security analysis tools (Semgrep, ease; Weggli-RS, ease). Weggli is a robust and efficient semantic code analysis for C and C++ programs, and Semgrep is designed for programs written in languages such as C#, Java, Java, JavaScript, Python, PHP, Ruby, Scala, etc.
LLMSecGuard relies on CyberSecEval, a benchmark specifically designed to evaluate the cybersecurity aspects of LLMs functioning as coding assistants (Bhatt et al., 2023). It comprises exhaustive test cases to assess the potential of LLMs in producing insecure code and facilitating cyberattacks. Nonetheless, users have the flexibility to update this benchmark or replace it with one of their preferences.
LLMSecGuard includes a few execution parameters as well. The primary one is the “termination condition” designed to prevent the system from looping indefinitely. This condition determines how many iterations a code piece should undergo improvement and analysis if potential vulnerabilities persist.
The “benchmark interval” parameter determines how often the benchmark process is executed, with options such as biweekly or monthly intervals. Given that LLMs evolve and their security properties may change, regular updates are necessary to ensure that results remain current.
Lastly, users can choose the LLMs of interest for benchmarking, select specific code analysis engines, and specify challenges that should be considered for measurement.
4.2. Security Benchmarking
In this scenario, a set of existing prompts, referred to as benchmark prompts, undergo evaluation. In particular, the “Benchmark Agent” sends each prompt in the benchmark to the “Prompt Agent”. This component dispatches the prompt to every LLM considered for benchmarking, collects responses for each prompt, and forwards them to the “Benchmark Agent”. Each prompt presents a challenge, and “Benchmark Agent” determines the extent to which an LLM succeeds in resolving it by comparing the response to the expected outcome. For instance, in terms of vulnerable code, it measures the quantity and severity of potential CWEs. Each LLM is assigned a performance score for each challenge, and ultimately, LLMs are ranked.
At present, LLMSecGuard relies on the CyberSecEval benchmark, which incorporates extensive prompts and regular expressions to uncover potential issues.
4.3. Secure Code Generation
Figure 2 illustrates the workflow in this scenario. It begins with the user providing a prompt of interest, anticipating the output to be a piece of code. The “Prompt Agent” assumes control and forwards the user’s prompt to the top LLM determined in the benchmarking scenario (unless the user chooses a different LLM). Upon receiving the LLM’s response, this component then transfers the produced code to the “Security Agent” for security inspection. The “Security Agent” collaborates with external analysis engines (configured earlier) to uncover potential vulnerabilities and respective lines in the code. If the termination condition is met (i.e., there is no vulnerability or the maximum analysis iterations is reached), the code along with its vulnerability information, is immediately reported back to the user. In cases where the termination condition is not met, this information is relayed to the “Prompt Agent”. This component formulates a new prompt based on the collected vulnerability information and queries the LLM once again.
Figure 2. LLMSecGuard’s secure code generation workflow
5. Relate Work
The popularity of AI-generated code has attracted the attention of security researchers. Pearce et al. assessed Copilot’s performance in suggesting code across 89 scenarios aligned with MITRE’s “Top 25” CWEs, revealing that approximately 40% of the generated programs contained vulnerabilities (Pearce et al., 2022). Nonetheless, in a later study, Asare et al. came to a different conclusion (Asare et al., 2023). Precisely, Copilot did not consistently reproduce past vulnerabilities introduced by humans. In about 25% of cases, Copilot even proposed fixes for these vulnerabilities, suggesting a potentially lower susceptibility to security risks compared to human programmers. Fu et al. analyzed 435 code snippets generated by Copilot in public GitHub projects and found that 35% of the snippets exhibit CWEs (Fu et al., 2023). Mousavi et al. designed 48 programming tasks for five common security APIs in Java and evaluated the performance of ChatGPT in developing these tasks. They uncovered a concerning level of security, specifically, they found an average misuse rate of 70% in these tasks (Mousavi et al., 2024).
Researchers have also developed benchmarks for assessing the security of LLM-generated code. Bhatt et al. developed CYBERSECEVAL, a benchmark tailored to evaluate the cybersecurity risks posed by LLMs (Bhatt et al., 2023). Hajipour et al. examined code models for generating high-risk security weaknesses and built a collection of diverse non-secure prompts for various vulnerability scenarios, which can be used as a benchmark to compare security weaknesses in LLMs (Hajipour et al., 2024).
6. Future Work
We plan to investigate whether LLMSecGuard will effectively support developers in writing secure code in real-world scenarios. Two groups of developers, both utilizing LLMs for coding, will be recruited for our study. We will assign the same programming tasks to each group, instructing one to utilize LLMs freely and the other to exclusively use LLMSecGuard during the coding process. We will measure the time taken to complete each task, the number, and the severity of vulnerabilities. We will compare the results across two groups as well as based on participants’ experience.
Pending positive evaluations, our ultimate goal is to integrate LLMSecGuard into at least one popular IDE (Integrated Development Environment), as deemed essential by developers for a seamless user experience (Tymchuk et al., 2018). Furthermore, this integration would allow LLMSecGuard to gather development context, thereby enabling the formulation of better prompts for code generation.
Future work could also investigate code changes in each iteration between “Prompt Agent” and “Security Agent”, as well as examine how the engineering of prompts and vulnerability information affects the performance of LLMs.
Finally, it is noteworthy that although a code snippet may not inherently contain vulnerabilities, its integration into a program or execution within a specific environment could still pose security risks and therefore warrants investigation (Stiévenart et al., 2022).
7. Conclusion
We introduced LLMSecGuard, an open-source framework developed to equip developers with code solutions that are more secure than the code initially generated by Large Language Models (LLMs). These code suggestions are obtained through the integration of LLMs and static security code analyzers. LLMSecGuard also measures the security properties of LLMs and provides researchers and practitioners with a current security comparison of different LLMs in the wild.

### v2

@@
-LLM Security Guard for Code
+Preprint Version
 Abstract.
@@
 Nevertheless, these models have exhibited limited capabilities in the security domain.
-We introduce LLMSecGuard, an open-source framework that offers enhanced code security through the synergy between static code analyzers and LLMs.
-LLMSecGuard aims to equip practitioners with code solutions that are more secure than the code initially generated by LLMs.
-It also benchmarks LLMs, providing valuable insights into the evolving security properties of these models.
+We introduce LLMSecGuard, a framework to offer enhanced code security through the synergy between static code analyzers and LLMs.
+LLMSecGuard is open source and aims to equip developers with code solutions that are more secure than the code initially generated by LLMs.
+This framework also has a benchmarking feature, aimed at providing insights into the evolving security attributes of these models.
 Security analysis, secure code generation, code models
-journalyear: 2024conference: 28th International Conference on Evaluation and Assessment in Software Engineering; June 18–21, 2024; Salerno, Italybooktitle: 28th International Conference on Evaluation and Assessment in Software Engineering (EASE 2024), June 18–21, 2024, Salerno, Italycopyright: nonedoi: 10.1145/xxxccs: Security and privacy Software security engineering
+journalyear: 2024copyright: rightsretainedconference: 28th International Conference on Evaluation and Assessment in Software Engineering; June 18–21, 2024; Salerno, Italybooktitle: 28th International Conference on Evaluation and Assessment in Software Engineering (EASE 2024), June 18–21, 2024, Salerno, Italydoi: 10.1145/3661167.3661263isbn: 979-8-4007-1701-7/24/06copyright: nonedoi: 10.1145/xxxccs: Security and privacy Software security engineering
 1. Introduction
@@
 Therefore, a vulnerable code piece that an LLM mistakenly recommends as a secure solution could compromise the entire system’s security if it is used without enough scrutiny.
-We introduce LLMSecGuard, a framework designed to extensively examine the security properties of LLMs and leverage their analytical capabilities for secure code development.
+We introduce LLMSecGuard, a framework specifically designed to examine the security properties of LLMs for code generation and harness their analytical capabilities to facilitate secure code development.
 It applies static security analysis on LLM-generated code to uncover potential security issues and guides LLMs in resolving such issues in the code.
-Moreover, LLMSecGuard assesses the security properties of LLMs and benchmark them across different CWEs.
+Moreover, LLMSecGuard can assess the security properties of LLMs and benchmark them across different CWEs.
 In summary, LLMSecGuard contributes to more secure software development.
@@
 The suggested code, at line 23, included the “SHA-256” hash function for key generation, which is not a best practice.
-Anyway, when we asked about the security risk at line 6 (i.e., static salt), ChatGPT responded: “Using a hard-coded string as a salt is generally not considered a good practice for security.
+Anyway, when we asked about the security risk at line 6 (ie., static salt), ChatGPT responded: “Using a hard-coded string as a salt is generally not considered a good practice for security.
 Ideally, salt should be unique for each user and should be randomly generated for each password.
@@
 3. LLMSecGuard
-We introduce LLMSecGuard, an open-source framework that offers enhanced code security through the synergy between code security analyzers and LLMs.
+We introduce LLMSecGuard, an open-source framework to offer enhanced code security through the synergy between code security analyzers and LLMs.
 The primary objective of LLMSecGuard is (i) to enhance the security of LLM-generated code, and (ii) to benchmark the security properties of LLMs.
 It adopts a RESTful architecture, implemented in Python using Django and Flask, and uses SQLite for data persistence.
-LLMSecGuard is publicly available on GitHub.
-[^3] https://github.com/aryakvnust/LLMSecGuard
+Figure 1 illustrates the schematic architecture of LLMSecGuard.
+We explain the main components in the following.
 Figure 1. LLMSecGuard’s components
-Figure 1 illustrates the schematic architecture of LLMSecGuard, comprising three main components: “Prompt Agent”, “Security Agent”, and “Benchmark Agent”.
 Prompt Agent.
@@
 This component puts different LLMs to security test.
-Particularly, it evaluates the security properties of LLMs based on a set of challenges, structured in JSON format.
+Particularly, it evaluates the security properties of LLMs based on a set of challenges structured in JSON format.
 Each challenge includes a prompt and the expected outcome, and “Benchmark Agent” is responsible for flagging LLMs that pass the challenge.
@@
 The “Security Agent” collaborates with external analysis engines (configured earlier) to uncover potential vulnerabilities and respective lines in the code.
-If the termination condition is met (i.e., there is no vulnerability or the maximum analysis iterations is reached), the code along with its vulnerability information, is immediately reported back to the user.
+If the termination condition is met (i.e., there is no vulnerability or the maximum analysis iterations is reached), the code, along with its vulnerability information, is immediately reported back to the user.
 In cases where the termination condition is not met, this information is relayed to the “Prompt Agent”.
@@
 These code suggestions are obtained through the integration of LLMs and static security code analyzers.
-LLMSecGuard also measures the security properties of LLMs and provides researchers and practitioners with a current security comparison of different LLMs in the wild.
+LLMSecGuard also offers a feature to measure the security properties of LLMs and provide a current security comparison of different LLMs in the wild.